The switch in an effective LR series dc-circuit
Answer #1

19)
L = 0.5 H
i = 9.5 A
di/dt = 1.1 A/s

V(t) = i*R + L*di/dt
V(t) = 9.5*R + 0.5*1.1
V(t) = 9.5*R + 0.55

At t = t
I(t) = Vs/R * ( 1 - e^-1)
9.5 = (9.5*R + 0.55)/R *  ( 1 - e^-1)
R = 0.1 Ω

Internal resistance, R = 0.1 Ω